OpenStudy (anonymous):

what is the multiplicative inverse of 4/5

OpenStudy (luigi0210):

Hello @childspaly Welcome to openstudy :)

OpenStudy (luigi0210):

And as for your question, what would you multiply the fraction by to get a 1?

OpenStudy (anonymous):

you just flip the fraction

OpenStudy (anonymous):

thanks but i have already been a meber hehehehehe

OpenStudy (anonymous):

hmmm, thats what i thought @Mad_e13

OpenStudy (ashleyisakitty):

(4/5*x=1)1÷(4/5)=1x5/4 = ?

OpenStudy (anonymous):

but @Luigi0210 is right too. if you flip the fraction and then multiply that by the original fraction you get 1

OpenStudy (jdoe0001):

hmmm the "multiplicative inverse" is just the fancy schmanzy name for the same expression flipped upside down